OpenStudy (anonymous):

solve the equation m+5(m-1)=7 (i know that m = 2 but i need steps plz!!!)

OpenStudy (anonymous):

\[m+5(m-1) = 7\] \[m+5m-5 = 7\] \[6m = 12\] \[m = 12/6\] m=2